### v2

# A Gated Residual Kolmogorov-Arnold Networks for Mixtures of Experts

Abstract
This paper introduces KAMoE, a novel Mixture of Experts (MoE) framework based on Gated Residual Kolmogorov-Arnold Networks (GRKAN). We propose GRKAN as an alternative to the traditional gating function, aiming to enhance efficiency and interpretability in MoE modeling. Through extensive experiments on digital asset markets and real estate valuation, we demonstrate that KAMoE consistently outperforms traditional MoE architectures across various tasks and model types. Our results show that GRKAN exhibits superior performance compared to standard Gating Residual Networks, particularly in LSTM-based models for sequential tasks. We also provide insights into the trade-offs between model complexity and performance gains in MoE and KAMoE architectures.
Figure 1: KAMoE
§§footnotetext: These authors contributed equally.
I Introduction
Introduced by [1, 2], the Mixture of Experts (MoE) is one of the most popular machine learning techniques. It involves the combination of multiple neural networks, where each network called “expert”, focusing on specific parts of the input. A gating network is used to assign different weights to these "experts". This gating mechanism is deciding which expert should be most trusted for a given input. This approach helps in improving the overall performance of the system by leveraging the strengths of individual neural networks specialized in different tasks. The MoE has since been applied to various domains such natural language processing [3, 4], computer vision [5, 6, 7], and reinforcement learning [8]. Time series data can exhibit complex, non-linear patterns and regime shifts [9] . The MoE’s ability to combine multiple specialized expert networks can help capture these complex temporal dynamics observed in financial time series [10]. Despite the increase in the availability of datasets for time series forecasting, it remains a very complex task. Powerful models have been proposed through the years [11, 12]. Some research proposed dynamic systems to model different existing states within a time series [13, 14, 9]. In this paper, we explore the Mixture of Experts (MoE) using a novel neural networks architecture embedded in a gating mechanism. This new gating mechanism is built over Kolmogorov-Arnold Networks [15], allowing for a more accurate estimation of weights for each neural network denoted $f_{i,\theta}\text{ and }i\in[0;n]$.
Recently, Liu et al. [15] released the Kolmogorov-Arnold Networks (KANs) a promising architecture introduced as an alternative for MLPs. To dive into the architecture of the model, we redirect the reader to [15]. In our previous work we have adapted KANs for time series forecasting [16], other researchers also proposed an adaptation for time series [17]. According to [18], KANs outperform conventional MLPs in real-world forecasting tasks. In a previous work we also proposed a temporal transformer architecture using TKANs [19]. Other extensions using wavelets [20] have been proposed in the meantime. TKANs aim to develop a framework with two key functionalities: handling sequential data and managing memory. To do so, in our previous work [16] we enriched the initial model by adding a recurring layer of KANs to introduce RKANs. Temporal Kolmogorov-Arnold Networks (TKANs) is an upgraded version of LSTM [21] using Kolmogorov-Arnold Networks (KANs). They rely on RKANs for the management of short-term memory as well as a cell state; for further details, we refer the reader to [16].
[…]
II Framework
In this paper, we introduce a new framework called “KAMoE” Figure 1, based on Gated Residual Kolmogorov-Arnold Networks (GRKAN) introduced in our previous work [22]. Each expert in charge of focusing on a specific part of the sequence requires optimal weighting. This weighting is achieved through a gating mechanism. A gated network is used to assign different weights to these “experts”. This gating mechanism is deciding which expert should be most trusted for a given input. This approach helps in improving the overall performance of the system by leveraging the strengths of individual neural networks specialized in different tasks. To assign these weights we will use the GRKAN.
II-A Inputs
The KAMoE is able to manage sequential and non-sequential data. It allows to leverage the experts according to specific tasks. We define $X_{i}=(X_{i,1},X_{i,2},...,X_{i,q})$ the vector of inputs where $X_{i,j}\in\mathbb{R}^{s},\quad j\in[1,2,...,q]$ and $s$ denotes the length of the input sequence. The particularity in the framework is that the input will be considered as learnable inputs denoted $\tilde{x}$ such
$\tilde{x_{i}}=w_{\tilde{x_{i}}}\odot X_{i},$ (1)
where $W_{\tilde{x_{i}}}\in\mathbb{R}^{dim(X)}$. The dimension of X noted $dim(X)$ may vary according to the inputs. In the case of sequential data we will have a 2D weight matrix. This transformed input with learnable parameter will be the input of the gated mechanism as well as all the experts defined in a latter section.
II-B Gated Mechanism
Establishing relationship between temporal data is a key issue. Gated Residual Networks (GRNs) offer an efficient and flexible way of modelling complex relationships in data. They allow to control the flow of information and facilitate the learning tasks. They are particularly useful in areas where nonlinear interactions and long-term dependencies are crucial. In our model we use the Gated Residual Kolmogorov-Arnold Networks (GRKAN) inspired from the GRN proposed by [23], we kept the same architecture. We propose a new approach using two KAN Linear layers [15] , to control the information flow while bringing more interpretability. Using GRKAN there is no more need for context which is contains in path signature, however, an additional linear layer is required to match the signature transform and the ouput of the gating mechanism.
Figure 2: Gated Residual KAN (GRKAN)
$\text{GRKAN}_{\omega}\left(x\right)$ $=\text{LayerNorm}\left(x+\text{GLU}_{\omega}(\eta_{1})\right),$ (2)
$\eta_{1}$ $=\text{KAN}(\varphi_{\eta_{1}}(.),\eta_{2}),$ (3)
$\eta_{2}$ $=\text{KAN}(\varphi_{\eta_{2}}(.),x).$ (4)
In this context, we used activation functions for KAN layers denoted, $\varphi_{\eta_{1}}(.)$ and $\varphi_{\eta_{2}}(.)$, SiLU [24] and ELU [25], respectively, while $\eta_{1}\in\mathbb{R}^{d_{model}}$ and $\eta_{2}\in\mathbb{R}^{d_{model}}$ represent intermediate layers. The standard layer normalization LayerNorm is that described in [26], and $\omega$ is an index used to indicate weight sharing. When the expression $\text{KAN}(x)$ is largely positive, ELU activation works as an identity function. On the other hand, when this expression is largely negative, ELU activation produces a constant output, thus behaving like a linear layer. We use Gated Linear Units (GLUs) [27] to provide the flexibility to suppress any parts of the architecture that are not required for a given dataset. Letting $\gamma\in\mathbb{R}^{d_{model}}$ be the input, the GLU then takes the form:
$\text{GLU}_{\omega}(\gamma)$ $=\sigma(W_{4,\omega}~{}\gamma+b_{4,\omega})\odot(W_{5,\omega}~{}\gamma+b_{5,\omega}),$ (5)
where $\sigma(.)$ is the sigmoid activation function, $W_{(.)}\in\mathbb{R}^{d_{model}\times d_{model}},b_{(.)}\in\mathbb{R}^{d_{model}}$ are the weights and biases, $\odot$ is the element-wise Hadamard product, and $d_{model}$ is the hidden state size. GLU allows to control the extent to which the GRN contributes to the original input $x$ – potentially skipping over the layer entirely if necessary as the GLU outputs could be all close to 0 in order to suppress the nonlinear contribution.
II-C Experts
We denote $f_{k,\theta_{k}}(.)$ a highly parametrized function which describes the succession of operations of neural networks such GRU, LSTM and TKAN. Let us denote $m$ the total number of experts, $k\in\{1,...,m\}$. The role of the expert in the framework is the learn from patterns within the sequence in the dataset. Each output of the expert is denoted $\tilde{y}_{k}$ and obtained simultaneously from the transformed input defined Eq. (1)
$\tilde{y}_{i,k}=f_{k,\theta_{k}}(\tilde{x_{i}}).$ (6)
All the estimates per experts will be stacked in $\tilde{y}_{i}$
$\tilde{y}_{i}:=\text{Concat}[\tilde{y}_{i,1},\tilde{y}_{i,2},...,\tilde{y}_{i,m}].$ (7)
II-D Output
The weight of each expert will be determined using the output of the GRKAN Eq.(4), for simplication we will denote the succession of the GRKAN operation $\Xi(.)$. Before the evaluation of these weights.
$z_{i}=\Psi_{\psi}(\tilde{x_{i}}).$ (8)
The output of this transformation will feed the GRKAN layer to obtain the weight. These outputs will be converted to weights,
$a_{i}=\sigma(\Xi(z_{i}))$ (9)
where $a_{i}\in\mathbb{R}^{m}$ and contains each weight for each expert and $\sigma(.)$ is a sigmoid activation function. The global output of our mixture will be obtained combining the weights and the ouput of each expert such
$\begin{split}\hat{y}_{i}&=\sum_{k}^{m}a_{i,k}f_{k,\theta_{k}}(x)\\ &=\sum_{k}^{m}a_{i,k}\tilde{y}_{i,k}\end{split}$ (10)
where $f_{k,\theta}(.)\in\mathbb{R}^{out}$ where $out$ is the output dimension of our mixture.
III Learning
As mixture of experts is a method that can be applied to any networks, we created two learning tasks in order to test it in both a sequential case with temporality and in a more standard one.
The first learning task is inspired by the one in [16], [19] and [22] with a few modifications. The task is about predicting market trading notional multiple steps ahead on one asset using the past values of the asset but also others. We chose this task as it’s a challenging one to perform due to the very noisy series but also the internal patterns in the traded volume like strong auto-correlation and seasonality.
We differ from the previous papers by reducing the number of input assets, which are BTC, ETH, ETC, XRP, BCH, LTC, and not only predicting BTC but instead training models and analyzing results for each of them. Everything else is similar to the previous papers; the data are from Binance and cover the period from January 1, 2020 to December 31, 2022. The data are preprocessed by dividing the values in the series by the moving median of the last two weeks and applying a MinMax Scaling between 0 and 1.
The model is kept simple to study the effect of MoE and KAMoE on them. As we are on a sequential prediction task, we used the two most standard RNN layers that are Gated Recurrent Unit (GRU), and Long Short Term Memory (LSTM). Each model is done by applying a first layer that returns the full sequence, followed by a second that only returns the last element. Finally, this output is fully connected with a linear layer that has the same number of units as there are steps to predict. The two hidden RNN layers are both composed of 100 units. Finally, to test the MoE and KAMoE effect, we only apply it on the first hidden RNN layer, which is composed of 3 experts and uses a sigmoid activation on the GRKAN outputs.
The second learning task is a more standard one, utilizing the California Housing dataset, which is commonly used in ML teaching and research. We obtained this dataset from scikit-learn. This dataset, derived from the 1990 U.S. census, focuses on housing in California and contains 20,640 instances with 8 numeric features. These features include median income, house age, average number of rooms, population, and geographical coordinates (latitude and longitude). The target variable for this task is the median house value for California districts, expressed in hundreds of thousands of dollars. This presents a regression problem, where the goal is to predict the median house value based on the given features.
For preprocessing the California Housing data, we applied StandardScaler fitted on the training set. The StandardScaler standardizes features by removing the mean and scaling to unit variance, which is indeed a good and standard approach for many machine learning algorithms, especially when dealing with features on different scales. This preprocessing step helps to ensure that all features contribute equally to the model and can improve the convergence of many machine learning algorithms.
We chose this dataset for several reasons:
  1. 1.
    It is not a sequential one so we can test the usefulness of our KAMoE and MoE in general not on time series
  2. 2.
    It represents a more traditional machine learning task, contrasting with our first, time-series-based task.
  3. 3.
    Its medium size and real-world nature provide a good balance between complexity and practicality.
  4. 4.
    As a widely-used benchmark, it facilitates comparison with other machine learning approaches in the literature.
By applying our mixture of experts method to these two distinct learning tasks, we aim to demonstrate its versatility across different types of data structures and problem domains.
For this task we compare the KAMoE effect on non-sequential model, we thus used very simple model which are an MLP and a KAN model. The MLP is composed of successive layers with ReLU activation, while not applying other activation to the KAN. We wanted to take the opportunity of this task to see the effect of the number of layers and hidden size for both, as well as being able to analyze whether the addition of the mixture framework does not increase results just by increasing the number of units. Finally we also compared the results with the one obtained with standard machine learning model using the one from scikit-learn as well as XGBoost.
III-A Task 1: Results
The results from the first task reveal that Mixture of Experts (MoE) is not a universally applicable solution for improving simple Recurrent Neural Networks (RNNs). Instead, its effectiveness varies depending on the specific context and model architecture. This variability is evident in two key aspects:
  1. 1.
    Differential impact on GRU vs. LSTM: The MoE framework shows markedly different effects when applied to Gated Recurrent Units (GRUs) compared to Long Short-Term Memory (LSTM) networks. While it generally yields significant improvements for LSTM models, the results for GRU-based models are more nuanced and context-dependent.
  2. 2.
    Asset-specific performance: The effectiveness of MoE varies across different cryptocurrencies. For instance, while it demonstrates notable improvements in Bitcoin (BTC) predictions, its performance on other assets is less consistent and more debatable.
These findings underscore that MoE is a sophisticated framework that should be applied judiciously, considering the specific characteristics of the problem at hand and the base model architecture.
A significant outcome of this study is the performance of our novel Kolmogorov-Arnold gating residual network (GRKAN), which replaces the standard Gating Residual Network (GRN) in the MoE framework. The GRKAN consistently outperforms the standard GRN across most scenarios, demonstrating the efficacy of incorporating Kolmogorov-Arnold principles into the gating mechanism of the mixture.
Examining the global average differences versus standard models, we observe:
  • For GRU-based models:
    • MoE shows a slight average decrease in performance (-0.007)
    • KAMoE demonstrates a small average improvement (0.005)
  • For LSTM-based models:
    • Both MoE and KAMoE show more substantial improvements (0.015 and 0.017 respectively)
These results highlight that while MoE and KAMoE architectures can offer improvements, their effectiveness is not uniform across all scenarios. The KAMoE variant, in particular, shows promise by consistently outperforming or matching the standard MoE in most cases.
III-B Task 2: Results
TABLE I: KANLinear: Mean R2 Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.767 0.762 0.766 0.755 0.758 0.769 0.757 0.751 0.621 0.691 0.688 0.671
10 0.765 0.777 0.760 0.753 0.768 0.771 0.758 0.751 0.621 0.705 0.697 0.683
25 0.768 0.770 0.753 0.744 0.772 0.767 0.750 0.736 0.621 0.699 0.694 0.682
50 0.776 0.765 0.751 0.745 0.775 0.761 0.745 0.747 0.622 0.699 0.686 0.684
100 0.783 0.765 0.760 0.756 0.780 0.762 0.748 0.758 0.621 0.693 0.684 0.685
200 0.783 0.762 0.765 0.770 0.784 0.763 0.768 0.765 0.622 0.693 0.689 0.698
400 0.781 0.748 0.768 0.772 0.785 0.746 0.768 0.771 0.622 0.689 0.699 0.705
800 0.782 0.741 0.765 0.773 0.789 0.742 0.769 0.773 0.622 0.696 0.703 0.708
Avg. Diff. vs. Standard 0.154145 0.065672 0.068344 0.068918 0.154954 0.064504 0.065217 0.067085 - - - -
The results obtained from our experiments yield several significant insights into the performance of various models on the given task.
III-C Comparison of KAN and MLP Models
Our findings indicate that the KAN model, when used independently, falls short of matching the MLP (Multi-Layer Perceptron) in terms of efficiency. However, both models exhibit similar patterns when increasing the number of units or hidden layers. Specifically:
  1. 1.
    Increasing the number of units generally improves model performance, albeit at the cost of increased complexity.
  2. 2.
    The optimal number of hidden layers appears to be two for both models, with additional layers providing negligible benefits.
Model R2 RMSE
KNeighborsRegressor 0.67001 0.432422
Lasso -0.000219 1.310696
LinearRegression 0.575788 0.555892
RandomForest 0.774564 0.295413
Ridge 0.575816 0.555855
SVR 0.727564 0.357003
XGBoost 0.537515 0.606044
TABLE II: Comparison of Machine Learning Models
III-D Impact of Mixture of Experts (MoE) and Kernel Activation Mixture of Experts (KAMoE)
The introduction of MoE and KAMoE architectures demonstrates substantial improvements in performance, particularly for smaller models. However, this effect diminishes for MLPs with a high number of units. In contrast, for the KAN model, the adoption of mixture architectures proves to be a game-changer, consistently enhancing performance across various configurations. Notably:
  • For KAN models, the ideal number of layers becomes one when using mixture architectures, reducing model complexity.
  • The number of parameters increases significantly with the number of units in MoE and KAMoE models, due to our choice of using an equal number of hidden units across all layers. This approach may not be optimal and could be refined to reduce the parameter count.
III-E Performance of KAMoE
One of the most intriguing findings is that KAMoE consistently outperforms MoE for both model types. While the standalone performance of KAN may be debatable, it undeniably serves as a valuable tool for enhancing model performance when strategically implemented.
III-F Comparison with Standard Machine Learning Models
We also evaluated standard machine learning models for comparison:
  • Neural network approaches generally yielded superior results compared to traditional machine learning models.
  • The Random Forest model stood out among traditional methods, achieving performance comparable to neural networks. However, it’s worth noting that Random Forest models can be highly complex, potentially rivaling neural networks in this aspect.
TABLE III: MLP: Mean R2 Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.753 0.764 0.769 0.768 0.755 0.762 0.767 0.758 0.695 0.728 0.726 0.727
10 0.768 0.775 0.772 0.774 0.764 0.769 0.776 0.780 0.733 0.755 0.759 0.764
25 0.775 0.777 0.782 0.783 0.775 0.783 0.781 0.783 0.751 0.773 0.780 0.785
50 0.777 0.786 0.788 0.787 0.774 0.785 0.787 0.789 0.756 0.781 0.785 0.786
100 0.781 0.791 0.795 0.792 0.782 0.788 0.792 0.792 0.762 0.790 0.794 0.793
200 0.783 0.792 0.796 0.797 0.784 0.792 0.794 0.797 0.764 0.794 0.795 0.796
400 0.785 0.796 0.799 0.797 0.786 0.798 0.798 0.795 0.765 0.801 0.800 0.797
800 0.788 0.798 0.801 0.798 0.786 0.797 0.798 0.798 0.772 0.803 0.800 0.800
Avg. Diff. vs. Standard 0.027 0.007 0.0.008 0.006 0.026 0.006 0.007 0.005 - - - -
IV Conclusion
In this paper, we introduced KAMoE, a novel framework for Mixture of Experts (MoE) modeling based on Gated Residual Kolmogorov-Arnold Networks (GRKAN). Our comprehensive experiments across both sequential and non-sequential tasks yield several significant insights:
  1. 1.
    Context-Dependent Efficacy: While MoE and KAMoE demonstrate potential for improving model performance, their effectiveness varies depending on the specific context and model architecture. This underscores the importance of judicious application of these techniques.
  2. 2.
    GRKAN Superiority: Our proposed GRKAN consistently outperforms standard Gating Residual Networks (GRN) across most scenarios, showcasing the efficacy of incorporating Kolmogorov-Arnold principles into the gating mechanism of MoE.
  3. 3.
    Model-Specific Impact: The impact of MoE and KAMoE varies across different model types. For instance, LSTM-based models showed more substantial improvements compared to GRU-based models when integrated with these techniques.
  4. 4.
    KAMoE Performance: KAMoE consistently outperformed or matched standard MoE in most cases, highlighting its potential as a powerful enhancement to existing MoE architectures.
  5. 5.
    Complexity vs. Performance Trade-off: While MoE and KAMoE architectures can offer significant performance improvements, they also increase model complexity and computational requirements. This trade-off needs careful consideration in practical applications.
  6. 6.
    KAN as a Strategic Tool: Although standalone Kolmogorov-Arnold Networks (KAN) may not always outperform traditional architectures like MLPs, our results demonstrate that KAN can be a valuable tool when strategically implemented within larger frameworks like KAMoE.
These findings open up new avenues for research in adaptive neural network architectures. Future work could explore optimizing the balance between model complexity and performance in KAMoE, investigating its applicability in other domains, and further refining the integration of Kolmogorov-Arnold principles in neural network design. Our work contributes to the ongoing effort to develop more efficient and adaptable machine learning models, potentially impacting a wide range of applications from time series forecasting to complex regression tasks.
Appendix A Additional Results Task 1
TABLE IV: Mean R2 Scores
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 0.398 0.391 0.392 0.399 0.371 0.385
3 0.239 0.227 0.233 0.129 0.178 0.205
6 0.090 0.115 0.155 -0.232 0.000 0.088
9 -0.045 0.073 0.054 -0.206 -0.047 -0.107
12 -0.117 -0.044 0.019 -0.449 -0.126 -0.080
15 -0.032 0.028 -0.017 -0.222 -0.142 -0.115
Avg. Diff. vs. Standard - 0.042 0.051 - 0.136 0.092
LTC 1 0.372 0.379 0.384 0.375 0.370 0.376
3 0.272 0.262 0.276 0.263 0.258 0.268
6 0.227 0.216 0.226 0.213 0.203 0.211
9 0.211 0.189 0.205 0.194 0.179 0.184
12 0.199 0.185 0.193 0.187 0.172 0.184
15 0.181 0.182 0.181 0.178 0.164 0.162
Avg. Diff. vs. Standard - -0.008 0.001 - -0.010 0.001
ETC 1 0.491 0.491 0.493 0.497 0.508 0.503
3 0.445 0.450 0.447 0.448 0.454 0.448
6 0.428 0.422 0.426 0.421 0.418 0.424
9 0.421 0.373 0.420 0.403 0.413 0.410
12 0.410 0.363 0.402 0.403 0.359 0.391
15 0.404 0.356 0.405 0.391 0.378 0.394
Avg. Diff. vs. Standard - -0.0240 -0.001 - -0.006 0.004
ETH 1 0.315 0.318 0.317 0.322 0.322 0.325
3 0.226 0.220 0.221 0.230 0.221 0.214
6 0.169 0.132 0.160 0.149 0.161 0.149
9 0.140 0.124 0.134 0.125 0.131 0.140
12 0.129 0.124 0.137 0.119 0.116 0.127
15 0.129 0.114 0.094 0.128 0.111 0.115
Avg. Diff. vs. Standard - -0.013 -0.007 - -0.002 +0.000
XRP 1 0.423 0.433 0.431 0.437 0.432 0.438
3 0.337 0.322 0.336 0.323 0.313 0.342
6 0.280 0.230 0.284 0.264 0.251 0.260
9 0.265 0.207 0.238 0.254 0.224 0.239
12 0.258 0.208 0.224 0.238 0.213 0.224
15 0.238 0.182 0.210 0.235 0.184 0.190
Avg. Diff. vs. Standard - -0.036 -0.013 - -0.022 0.0013
BCH 1 0.265 0.270 0.268 0.271 0.271 0.273
3 0.180 0.184 0.181 0.180 0.182 0.180
6 0.139 0.134 0.139 0.138 0.135 0.137
9 0.113 0.107 0.104 0.111 0.110 0.111
12 0.101 0.095 0.094 0.095 0.084 0.096
15 0.096 0.075 0.088 0.096 0.077 0.088
Avg. Diff. vs. Standard - -0.005 -0.003 - -0.005 0.002
Global Avg. Diff. vs. Standard - -0.007 0.005 - 0.015 0.017
TABLE V: Standard Deviation of R2 Scores
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 0.004 0.003 0.005 0.007 0.004 0.012
3 0.007 0.005 0.009 0.051 0.060 0.026
6 0.047 0.030 0.020 0.065 0.112 0.047
9 0.058 0.034 0.052 0.070 0.118 0.065
12 0.104 0.056 0.046 0.201 0.112 0.098
15 0.055 0.043 0.056 0.188 0.069 0.100
LTC 1 0.011 0.009 0.006 0.007 0.017 0.013
3 0.012 0.007 0.006 0.011 0.008 0.006
6 0.011 0.007 0.006 0.003 0.008 0.006
9 0.008 0.015 0.007 0.004 0.008 0.011
12 0.004 0.010 0.005 0.004 0.004 0.004
15 0.006 0.007 0.006 0.003 0.002 0.009
ETC 1 0.004 0.006 0.004 0.004 0.003 0.003
3 0.003 0.003 0.006 0.003 0.004 0.004
6 0.002 0.005 0.004 0.006 0.009 0.007
9 0.001 0.017 0.003 0.010 0.006 0.009
12 0.004 0.009 0.014 0.003 0.020 0.013
15 0.004 0.030 0.004 0.007 0.014 0.009
ETH 1 0.004 0.003 0.001 0.001 0.001 0.001
3 0.002 0.003 0.004 0.002 0.004 0.007
6 0.006 0.011 0.015 0.017 0.007 0.027
9 0.012 0.011 0.017 0.014 0.009 0.008
12 0.006 0.007 0.003 0.011 0.023 0.015
15 0.001 0.011 0.036 0.005 0.012 0.013
XRP 1 0.023 0.026 0.026 0.010 0.016 0.024
3 0.019 0.021 0.016 0.011 0.033 0.005
6 0.015 0.033 0.030 0.020 0.026 0.027
9 0.008 0.016 0.019 0.011 0.020 0.015
12 0.008 0.027 0.010 0.018 0.018 0.013
15 0.028 0.035 0.021 0.015 0.029 0.024
BCH 1 0.002 0.002 0.004 0.004 0.005 0.002
3 0.002 0.003 0.003 0.002 0.004 0.006
6 0.004 0.006 0.005 0.002 0.002 0.005
9 0.005 0.006 0.015 0.006 0.003 0.005
12 0.006 0.008 0.005 0.005 0.009 0.009
15 0.004 0.008 0.005 0.005 0.009 0.010
TABLE VI: Mean RMSE
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 0.055 0.055 0.055 0.055 0.056 0.056
3 0.062 0.062 0.062 0.066 0.064 0.063
6 0.068 0.067 0.065 0.078 0.071 0.068
9 0.072 0.068 0.069 0.077 0.072 0.074
12 0.074 0.072 0.070 0.085 0.075 0.073
15 0.072 0.070 0.072 0.078 0.076 0.075
LTC 1 0.015 0.015 0.015 0.015 0.016 0.015
3 0.017 0.017 0.017 0.017 0.017 0.017
6 0.017 0.018 0.017 0.018 0.018 0.018
9 0.018 0.018 0.018 0.018 0.018 0.018
12 0.018 0.018 0.018 0.018 0.018 0.018
15 0.018 0.018 0.018 0.018 0.018 0.018
ETC 1 0.030 0.030 0.030 0.030 0.030 0.030
3 0.032 0.032 0.032 0.032 0.032 0.032
6 0.033 0.033 0.033 0.033 0.033 0.033
9 0.034 0.035 0.034 0.034 0.034 0.034
12 0.034 0.036 0.035 0.034 0.036 0.035
15 0.035 0.036 0.035 0.035 0.035 0.035
ETH 1 0.053 0.053 0.053 0.053 0.053 0.053
3 0.057 0.057 0.057 0.057 0.057 0.057
6 0.059 0.060 0.059 0.059 0.059 0.059
9 0.059 0.060 0.059 0.060 0.060 0.059
12 0.059 0.059 0.059 0.059 0.059 0.059
15 0.059 0.059 0.060 0.059 0.059 0.059
XRP 1 0.009 0.009 0.009 0.009 0.009 0.009
3 0.009 0.009 0.009 0.009 0.009 0.009
6 0.010 0.010 0.010 0.010 0.010 0.010
9 0.010 0.010 0.010 0.010 0.010 0.010
12 0.010 0.010 0.010 0.010 0.010 0.010
15 0.010 0.011 0.010 0.010 0.011 0.011
BCH 1 0.036 0.036 0.036 0.036 0.036 0.036
3 0.037 0.037 0.037 0.037 0.037 0.037
6 0.038 0.038 0.038 0.038 0.038 0.038
9 0.039 0.039 0.039 0.039 0.039 0.039
12 0.039 0.039 0.039 0.039 0.039 0.039
15 0.039 0.039 0.039 0.039 0.039 0.039
TABLE VII: Standard Deviation of RMSE (×$\times 10^{-3}$)
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 0.163 0.129 0.213 0.323 0.161 0.521
3 0.283 0.208 0.347 1.877 2.232 1.021
6 1.674 1.088 0.746 2.104 3.837 1.684
9 1.977 1.204 1.835 2.163 3.884 2.135
12 3.377 1.920 1.616 5.708 3.550 3.239
15 1.907 1.523 1.939 5.881 2.232 3.223
LTC 1 0.138 0.107 0.073 0.084 0.210 0.157
3 0.140 0.085 0.070 0.127 0.086 0.065
6 0.122 0.076 0.062 0.032 0.084 0.066
9 0.086 0.162 0.080 0.040 0.087 0.124
12 0.043 0.106 0.053 0.045 0.049 0.042
15 0.063 0.074 0.065 0.036 0.026 0.093
ETC 1 0.112 0.170 0.131 0.108 0.095 0.104
3 0.094 0.091 0.162 0.085 0.112 0.122
6 0.069 0.145 0.114 0.169 0.258 0.200
9 0.027 0.461 0.081 0.294 0.164 0.242
12 0.118 0.248 0.407 0.102 0.551 0.365
15 0.118 0.813 0.111 0.203 0.381 0.259
ETH 1 0.139 0.122 0.052 0.038 0.053 0.027
3 0.090 0.109 0.137 0.060 0.126 0.239
6 0.212 0.386 0.529 0.566 0.252 0.910
9 0.394 0.383 0.566 0.474 0.306 0.265
12 0.198 0.237 0.085 0.364 0.758 0.506
15 0.037 0.355 1.166 0.160 0.400 0.419
XRP 1 0.175 0.195 0.194 0.076 0.120 0.182
3 0.127 0.144 0.107 0.075 0.225 0.033
6 0.098 0.214 0.202 0.135 0.173 0.180
9 0.057 0.099 0.122 0.074 0.131 0.100
12 0.056 0.174 0.071 0.118 0.118 0.087
15 0.184 0.221 0.133 0.102 0.190 0.151
BCH 1 0.059 0.055 0.089 0.087 0.133 0.050
3 0.052 0.070 0.067 0.048 0.099 0.142
6 0.098 0.136 0.118 0.054 0.051 0.101
9 0.098 0.122 0.317 0.134 0.067 0.101
12 0.129 0.168 0.095 0.115 0.181 0.181
15 0.080 0.169 0.103 0.102 0.177 0.211
TABLE VIII: Mean Training Time (seconds)
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 27.55 55.51 71.71 20.79 51.98 55.39
3 34.80 70.15 77.08 23.78 55.29 59.36
6 38.62 68.10 74.61 23.16 58.57 59.50
9 35.22 64.98 72.55 22.97 57.84 62.81
12 44.37 86.27 85.02 27.93 71.99 71.55
15 45.77 93.70 94.44 28.76 76.41 90.06
LTC 1 20.44 50.54 54.58 17.53 48.08 52.44
3 21.36 52.24 54.28 18.31 47.96 52.89
6 21.64 51.62 54.67 18.17 47.89 54.97
9 21.40 52.53 54.23 17.14 49.62 51.50
12 26.99 62.88 63.69 21.40 56.58 61.42
15 30.88 70.07 72.75 24.46 65.21 68.95
ETC 1 24.06 55.19 60.91 17.58 51.35 56.75
3 25.42 55.73 65.01 19.40 49.78 55.99
6 24.08 59.57 63.16 22.17 51.09 59.20
9 23.45 55.24 61.17 21.52 52.80 59.61
12 27.65 62.03 68.12 25.22 64.16 70.16
15 36.36 74.78 78.54 28.27 68.30 73.43
ETH 1 24.34 61.50 59.85 21.75 54.38 57.76
3 30.07 63.31 65.02 23.84 57.06 70.27
6 27.42 83.99 71.75 28.45 60.38 72.27
9 29.85 84.33 69.60 25.02 60.18 61.39
12 38.23 82.89 70.96 29.73 69.93 67.31
15 48.82 102.10 116.28 32.61 75.71 86.59
XRP 1 21.94 51.75 55.01 18.19 47.65 53.61
3 21.45 53.40 55.08 18.32 48.91 53.82
6 21.98 53.09 55.63 17.80 49.39 53.00
9 21.29 52.16 55.32 18.26 48.75 52.32
12 26.66 59.89 63.57 22.08 58.58 61.34
15 30.59 73.11 74.55 24.61 66.44 69.45
BCH 1 21.67 59.54 63.97 17.39 51.32 54.11
3 25.39 57.43 62.10 18.76 55.65 56.19
6 23.00 53.71 57.87 19.01 53.66 54.61
9 21.06 54.96 60.05 18.37 50.70 54.64
12 26.37 63.25 64.94 22.00 60.30 62.66
15 31.90 73.16 77.60 24.98 66.92 70.82
TABLE IX: Standard Deviation of Training Time (seconds)
Asset Steps GRU MoE_GRU KAMoE_GRU LSTM MoE_LSTM KAMoE_LSTM
BTC 1 3.87 3.01 7.41 2.04 2.80 5.13
3 4.00 4.10 9.39 2.27 6.01 3.46
6 3.93 9.09 10.20 2.24 5.40 5.83
9 3.51 4.04 7.66 1.37 6.53 7.51
12 8.86 14.73 9.19 2.71 10.77 5.45
15 4.79 13.13 10.39 3.11 6.39 11.62
LTC 1 0.49 1.24 2.19 1.13 2.42 2.66
3 1.29 3.65 2.32 0.79 1.75 2.58
6 0.64 2.22 1.55 0.98 1.22 4.37
9 1.57 3.66 1.81 0.13 3.00 1.89
12 3.79 4.64 1.64 1.73 1.81 2.69
15 1.04 3.36 2.02 0.56 2.21 1.74
ETC 1 2.99 3.56 7.21 1.26 2.42 5.91
3 3.63 4.00 8.66 1.64 2.69 3.96
6 2.09 2.05 6.27 1.89 4.80 5.56
9 2.30 2.33 3.43 3.10 5.86 6.33
12 1.97 0.84 4.93 3.25 6.04 6.78
15 4.41 5.73 6.32 3.06 5.02 3.68
ETH 1 3.94 4.38 6.62 2.14 3.45 4.08
3 4.35 8.34 3.13 1.79 6.15 6.20
6 4.71 5.53 5.09 2.81 2.92 7.43
9 5.96 9.74 9.77 2.38 3.64 4.85
12 6.33 7.85 5.48 3.50 6.85 3.59
15 3.69 11.27 12.16 2.77 5.49 5.87
XRP 1 1.76 1.81 2.27 1.13 1.74 2.12
3 0.29 2.91 2.21 1.11 1.72 4.38
6 1.06 3.98 2.01 0.52 2.77 3.53
9 0.55 1.11 2.16 1.43 1.33 2.47
12 1.60 1.22 1.54 0.85 4.50 2.06
15 1.10 8.40 2.15 0.41 3.42 0.58
BCH 1 0.61 5.45 7.53 1.32 2.00 3.05
3 4.17 2.12 3.64 1.66 7.21 3.89
6 1.70 2.52 3.11 1.79 3.77 3.22
9 0.33 3.34 3.05 1.54 2.89 3.32
12 1.07 2.35 2.15 0.88 2.76 1.32
15 1.34 6.48 4.19 1.10 3.90 2.45
Appendix B Additional Results Task 2
TABLE X: KANLinear: Standard Deviation of R2 Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.0061 0.0081 0.0057 0.0105 0.0068 0.0034 0.0044 0.0095 0.0016 0.0119 0.0124 0.0124
10 0.0099 0.0054 0.0089 0.0055 0.0059 0.0031 0.0106 0.0031 0.0007 0.0067 0.0050 0.0100
25 0.0039 0.0049 0.0074 0.0079 0.0033 0.0053 0.0045 0.0090 0.0008 0.0035 0.0022 0.0047
50 0.0062 0.0033 0.0085 0.0113 0.0040 0.0020 0.0071 0.0083 0.0007 0.0042 0.0089 0.0058
100 0.0043 0.0028 0.0057 0.0072 0.0030 0.0023 0.0123 0.0036 0.0007 0.0032 0.0040 0.0058
200 0.0027 0.0041 0.0038 0.0066 0.0029 0.0043 0.0056 0.0047 0.0004 0.0044 0.0040 0.0041
400 0.0033 0.0031 0.0058 0.0058 0.0031 0.0066 0.0039 0.0015 0.0004 0.0023 0.0041 0.0021
800 0.0049 0.0165 0.0180 0.0082 0.0055 0.0072 0.0029 0.0014 0.0001 0.0016 0.0020 0.0032
TABLE XI: MLP: Standard Deviation of R2 Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.0121 0.0126 0.0056 0.0053 0.0100 0.0079 0.0078 0.0090 0.0259 0.0135 0.0075 0.0075
10 0.0036 0.0048 0.0023 0.0053 0.0050 0.0079 0.0046 0.0045 0.0096 0.0053 0.0047 0.0041
25 0.0069 0.0043 0.0037 0.0033 0.0071 0.0043 0.0052 0.0035 0.0029 0.0024 0.0024 0.0037
50 0.0020 0.0033 0.0039 0.0021 0.0033 0.0022 0.0030 0.0023 0.0009 0.0029 0.0041 0.0039
100 0.0047 0.0056 0.0020 0.0037 0.0042 0.0021 0.0043 0.0029 0.0022 0.0007 0.0036 0.0011
200 0.0048 0.0034 0.0019 0.0023 0.0037 0.0044 0.0017 0.0025 0.0023 0.0032 0.0009 0.0032
400 0.0019 0.0028 0.0021 0.0022 0.0043 0.0038 0.0017 0.0037 0.0025 0.0015 0.0019 0.0016
800 0.0023 0.0033 0.0042 0.0023 0.0015 0.0008 0.0017 0.0020 0.0022 0.0011 0.0012 0.0025
TABLE XII: KANLinear: Mean MSE Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.305 0.312 0.307 0.321 0.317 0.303 0.319 0.326 0.497 0.405 0.408 0.432
10 0.308 0.292 0.314 0.324 0.304 0.300 0.317 0.327 0.496 0.387 0.396 0.416
25 0.304 0.302 0.323 0.335 0.299 0.306 0.327 0.346 0.496 0.395 0.401 0.417
50 0.294 0.308 0.327 0.334 0.294 0.313 0.335 0.332 0.496 0.394 0.411 0.414
100 0.285 0.308 0.315 0.320 0.288 0.312 0.330 0.317 0.496 0.402 0.414 0.413
200 0.284 0.311 0.308 0.302 0.283 0.310 0.304 0.308 0.496 0.403 0.408 0.396
400 0.286 0.330 0.304 0.299 0.281 0.333 0.304 0.300 0.496 0.408 0.395 0.386
800 0.286 0.340 0.308 0.298 0.277 0.338 0.303 0.297 0.495 0.399 0.389 0.383
TABLE XIII: MLP: Mean MSE Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.324 0.310 0.303 0.303 0.321 0.312 0.305 0.317 0.399 0.357 0.359 0.358
10 0.305 0.295 0.299 0.296 0.310 0.303 0.293 0.289 0.350 0.321 0.316 0.309
25 0.295 0.292 0.285 0.285 0.295 0.285 0.287 0.284 0.327 0.297 0.289 0.282
50 0.292 0.280 0.278 0.278 0.296 0.282 0.279 0.277 0.319 0.287 0.282 0.280
100 0.287 0.274 0.269 0.273 0.286 0.277 0.272 0.272 0.312 0.275 0.270 0.271
200 0.284 0.272 0.267 0.267 0.283 0.273 0.270 0.266 0.310 0.270 0.269 0.268
400 0.282 0.268 0.263 0.266 0.281 0.265 0.265 0.269 0.308 0.261 0.262 0.266
800 0.278 0.265 0.261 0.265 0.280 0.266 0.264 0.265 0.299 0.258 0.262 0.261
TABLE XIV: KANLinear: Standard Deviation of MSE Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.0080 0.0106 0.0075 0.0137 0.0089 0.0044 0.0057 0.0125 0.0021 0.0156 0.0162 0.0162
10 0.0129 0.0070 0.0117 0.0072 0.0077 0.0041 0.0139 0.0041 0.0009 0.0088 0.0066 0.0131
25 0.0051 0.0065 0.0097 0.0104 0.0044 0.0070 0.0059 0.0117 0.0010 0.0045 0.0029 0.0061
50 0.0081 0.0044 0.0111 0.0149 0.0053 0.0026 0.0092 0.0109 0.0010 0.0054 0.0117 0.0076
100 0.0056 0.0037 0.0074 0.0094 0.0039 0.0030 0.0161 0.0048 0.0010 0.0042 0.0052 0.0076
200 0.0035 0.0053 0.0049 0.0087 0.0039 0.0056 0.0073 0.0061 0.0005 0.0057 0.0052 0.0053
400 0.0043 0.0041 0.0076 0.0077 0.0040 0.0087 0.0051 0.0020 0.0005 0.0030 0.0054 0.0028
800 0.0064 0.0216 0.0236 0.0108 0.0072 0.0095 0.0038 0.0018 0.0001 0.0021 0.0026 0.0041
TABLE XV: MLP: Standard Deviation of MSE Values
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 0.0159 0.0166 0.0074 0.0069 0.0131 0.0104 0.0102 0.0118 0.0339 0.0177 0.0098 0.0098
10 0.0047 0.0062 0.0031 0.0069 0.0066 0.0104 0.0060 0.0060 0.0126 0.0069 0.0061 0.0054
25 0.0091 0.0056 0.0049 0.0043 0.0093 0.0056 0.0068 0.0046 0.0038 0.0031 0.0032 0.0049
50 0.0026 0.0043 0.0050 0.0027 0.0043 0.0029 0.0039 0.0030 0.0012 0.0038 0.0054 0.0051
100 0.0062 0.0073 0.0026 0.0048 0.0055 0.0027 0.0056 0.0038 0.0029 0.0009 0.0048 0.0014
200 0.0063 0.0045 0.0025 0.0030 0.0049 0.0058 0.0022 0.0033 0.0030 0.0042 0.0011 0.0042
400 0.0025 0.0037 0.0027 0.0029 0.0056 0.0050 0.0022 0.0049 0.0032 0.0019 0.0025 0.0021
800 0.0030 0.0044 0.0055 0.0030 0.0020 0.0010 0.0022 0.0027 0.0028 0.0015 0.0016 0.0032
TABLE XVI: KANLinear: Mean Time Values (in seconds)
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 55.32 100.64 118.63 138.04 66.54 84.89 96.58 108.80 14.45 23.73 20.42 38.63
10 56.62 67.07 83.68 103.41 47.33 66.07 74.86 85.02 10.39 17.15 16.92 19.61
25 41.72 54.45 65.45 89.20 32.87 45.56 52.63 72.48 9.14 12.54 12.36 14.83
50 44.43 50.70 65.12 83.17 33.07 42.58 53.72 67.16 7.73 10.46 11.12 13.38
100 40.57 49.21 61.75 80.74 33.08 38.73 54.64 68.04 7.95 10.38 11.24 13.79
200 45.69 48.19 65.67 82.30 34.24 42.10 53.80 71.12 9.04 10.34 11.73 14.67
400 43.84 50.22 70.00 89.24 37.25 42.66 60.32 72.29 8.39 10.34 12.64 15.71
800 47.18 54.07 82.58 100.57 33.36 48.52 69.21 79.60 8.99 10.49 15.02 18.15
TABLE XVII: MLP: Mean Time Values (in seconds)
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 50.61 79.67 75.23 80.75 28.42 43.67 53.50 67.08 17.59 22.59 18.95 14.87
10 36.79 44.95 72.77 84.54 43.09 41.17 65.27 71.60 17.18 20.15 16.13 22.19
25 35.26 42.62 59.26 75.78 29.39 37.58 40.08 53.66 21.15 21.98 15.06 15.33
50 30.36 37.90 55.01 65.47 20.66 34.10 38.92 48.91 19.03 16.77 12.98 15.20
100 29.63 45.48 55.32 69.46 23.34 33.94 39.38 45.23 16.73 15.48 10.36 10.16
200 24.52 47.69 55.19 62.29 24.27 30.05 39.53 46.21 14.10 19.50 12.22 10.16
400 28.09 46.29 52.46 65.45 25.57 34.68 35.86 44.78 12.58 15.60 12.64 12.32
800 38.75 53.89 52.69 59.20 22.93 32.79 39.06 42.71 11.56 14.13 9.55 8.93
TABLE XVIII: KANLinear: Mean Number of Parameters
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 1228 2099 2970 3841 1070 1783 2496 3209 317 517 717 917
10 2253 4949 7645 10341 2005 4453 6901 9349 607 1347 2087 2827
25 5328 20099 34870 49641 4810 19063 33316 47569 1477 5937 10397 14857
50 10453 67349 124245 181141 9485 65413 121341 177269 2927 20587 38247 55907
100 20703 244349 467995 691641 18835 240613 462391 684169 5827 76137 146447 216757
200 41203 928349 1815495 2702641 37535 921013 1804491 2687969 11627 292237 572847 853457
400 82203 3616349 7150495 10684641 74935 3601813 7128691 10655569 23227 1144437 2265647 3386857
800 164203 14272349 28380495 42488641 149735 14243413 28337091 42430769 46427 4528837 9011247 13493657
TABLE XIX: MLP: Mean Number of Parameters
Hidden Units KAMoE MoE Standard
Number of Layers
1 2 3 4 1 2 3 4 1 2 3 4
5 430 791 1152 1513 272 475 678 881 51 81 111 141
10 735 1541 2347 3153 487 1045 1603 2161 101 211 321 431
25 1650 4991 8332 11673 1132 3955 6778 9601 251 901 1551 2201
50 3175 14741 26307 37873 2207 12805 23403 34001 501 3051 5601 8151
100 6225 49241 92257 135273 4357 45505 86653 127801 1001 11101 21201 31301
200 12325 178241 344157 510073 8657 170905 333153 495401 2001 42201 82401 122601
400 24525 676241 1327957 1979673 17257 661705 1306153 1950601 4001 164401 324801 485201
800 48925 2632241 5215557 7798873 34457 2603305 5172153 7741001 8001 648801 1289601 1930401